# FocusChat: Text-guided Long Video Understanding via Spatiotemporal Information Filtering

### v1

@@ 1 Introduction
Figure 3: When asking a question about a five-minute video, such as ”How many people are playing basketball in the first two minutes? Describe it in detail,” the semantic similarity matrix H and the attention-map SAM diagram in STFM are presented.
@@ 2 Related Work
Vision Large Language Models (vLLMs). The integration of the perceptual abilities of the vision models [21, 35, 37, 40] with the reasoning capabilities of LLMs has given rise to Vision Large Language Models (vLLMs) [66, 17, 27]. VLLMs encompass both Image-Language and Video-Language models. This approach transforms visual signals into tokens interpretable by LLMs. Image-Language models integrate powerful pre-trained language models with image encoders to enhance multi-modal reasoning capabilities[23, 30, 73]. For instance, Flamingo [1] connects state-of-the-art vision-only and language-only models to excel in few-shot learning tasks. BLIP-2[23] introduces a lightweight querying transformer to bridge the gap between frozen image encoders and language models, achieving strong performance with fewer trainable parameters. LLaVA [30] utilizes a simple linear layer to project image features into the text embedding space, effectively fine-tuning language models for improved outcomes. MiniGPT-4 builds on BLIP-2 [73]by gathering a large dataset of image-text pairs, enhancing language generation. Video-language models have evolved from image-language models like Flamingo [1] and BLIP-2[23], which flatten spatio-temporal features into 1D sequences but struggle to capture temporal dynamics. Models such as Video-LLaMA[71] add video querying transformers to enhance temporal understanding, while Video-ChatGPT[33] averages spatial-level features for video representation. ChatVideo[53] uses tracklets annotated with textual descriptions and employs ChatGPT for various tasks, while VideoChat [24] generates action and object annotations for LLM reasoning.
@@
Long Video Understanding. Long video understanding poses significant challenges in computer vision, primarily because it requires capturing long-range patterns in videos that often exceed 30 seconds. A typical strategy[3, 55, 6] involves maintaining a memory bank to store historical information and utilizing parametric[55] or non-parametric[3] compression modules to manage this efficiently. Recent approaches [19, 16, 41] have also integrated language as a bridge for understanding, breaking long videos into shorter clips, generating textual descriptions for each, and then employing large language models (LLMs) to aggregate these captions for analysis. However, these methods are cumbersome and lack conciseness. Some have adopted streaming methods[16, 39, 72, 42] to process long videos. Although these methods appear promising, indiscriminately handling all video frames inevitably leads to information redundancy. [47] is a key-frame-based model, distinct from previous methods. Although it appears to treat each frame differently, the key frame is fixed and unrelated to the prompt. Our proposed text-guided approach effectively addresses this issue. There are very few video understanding models based on this method, with [39] being one of them. However, it has several drawbacks, including a complex model structure, time-consuming computations, and an inability to filter information at the frame level or even at the region level. In contrast, our FocusChat effectively overcomes these limitations.
@@ 3.1.1 Visual Semantic Branch
The visual semantic branch extracts the semantic information from each image. It consists of a frozen pre-trained image encoder and an image Q-Former. A video consists of $T$ frames is represented as $\mathbf{V}\in\mathbb{R}^{T\times H\times W\times 3}$. The input frames are passed into a pre-trained visual encoder, i.e., ViT, to obtain the visual frame features $\mathbf{V_{f}}\in\mathbb{R}^{T\times N_{v}\times C}$, where $N_{v}$ represents the number of patches, and $C$ denotes the number of feature channels. Subsequently, an image Q-Former further compresses the frame features. As Fig. 2 illustrates, the image Q-Former takes as input M learnable queries of length D. These queries interact with the frame features via cross-attention and update the initial queries to output the final M visual semantic vectors of length D, denoted as $\mathbf{V_{q}}\in\mathbb{R}^{T\times M\times D}$. Each visual vector contains semantic information at the region level or frame level. For example, the visual semantics extracted from the video in Fig. 1 may include elements such as men, women, vegetables, treadmills, etc.
@@ 3.1.2 Textual Semantic Branch
The text semantic branch is designed to semantically encode the user’s input prompt $P$. A pre-trained CLIP[40] model encodes the user prompt, resulting in prompt features $\mathbf{P_{f}}\in\mathbb{R}^{N_{p}\times D}$ , where $D$ denotes the dimension of text embedding, and $N_{p}$ is the text token number. We then employ a Q-Former, similar to the visual semantic branch, to encode $\mathbf{P_{f}}$. In this process, $\mathbf{P_{f}}$ is projected as keys and values. By utilizing M trainable query embeddings, we obtain M text vectors, denoted as $\mathbf{P_{q}}\in\mathbb{R}^{M\times D}$. Each textual vector contains semantic information at the word or subword level. For example, the prompt “How many people are cooking in the kitchen?” in Fig. 1 may be parsed into semantic vectors such as person, cooking, kitchen, etc.
@@ 3.2.1 Hybrid Spatial-Temporal Filtering Module
To obtain visual features that are semantically close to the user’s query, we first compute the semantic similarity matrix by calculating the similarity between text and visual semantic vectors. The similarity matrix is then used to guide the hybrid spatial-temporal filtering (HSTF) module for visual information filtering. The filtering process is explicitly applied to the video in both spatial and temporal dimensions. Spatial filtering is performed row by row, as shown in Fig. 2, while temporal filtering is applied column by column.
Semantic Similarity Matrix. We compute the cosine similarity between the text semantic feature $\mathbf{P_{q}}$ and the visual semantic feature $\mathbf{V_{q}}$ to get the similarity matrix $\mathbf{H}\in\mathbb{R}^{T\times M}$. As shown in Eq. 1, sim indicates cosine similarity and normalization, t is the frame index, and i is the semantic feature index. Since text semantics can achieve word-level granularity and visual semantics can achieve region-level granularity, therefore $\mathbf{H}$ represents fine-grained similarity. For example, in Fig. 3, each element of the semantic similarity matrix H represents a specific region in the image, such as an arm, body, or basketball. It serves as a guide for filtering the visual spatial-temporal features.
@@
Hybrid Spatial-Temporal Filtering Cross-Attention. The input to STFM consists of trainable queries, which are fed into self-attention, resulting in $\mathbf{V_{s}}\in\mathbb{R}^{N\times d}$, where N is the number of queries. Given $\mathbf{V_{s}}$, semantic similarity matrix $\mathbf{H}$, and the visual semantic feature $\mathbf{V_{q}}$, as shown in Fig. 2, we aim to produce an attention map that is closely related to the semantic similarity matrix. We first project $\mathbf{V_{q}}$ to generate keys and values using $\mathbf{W_{k}}$ and $\mathbf{W_{v}}$. Our hybrid spatial-temporal filtering module based attention-map $\mathbf{SAM}\in\mathbb{R}^{N\times TM}$ is given in Eq. 2, which is equivalent to $SAM=softmax(\alpha logH+QK^{T}/\sqrt{d})$, where $\alpha$ is a constant greater than or equal to 0. So $g(x)=\alpha logx$ as shown in Fig. 2. The output of the hybrid spatial-temporal filtering module is denoted as $\mathbf{Z}\in\mathbb{R}^{N\times d}$. Eq. 2 and Eq. 3 complete most of the spatiotemporal information filtering. Therefore, the redundant information filtered in $\mathbf{V_{q}}$ alleviates the pressure on the number of vision tokens sent to LLM.
@@
3.2.2 Prompt Based Temporal Filtering Module
The visual features highly relevant to the question are extracted with the help of HSTF. However, there are still some redundant temporal information unfiltered. For example, in the left image of Fig. 3, the semantic information from the last minute is encoded into the visual tokens. For these time-sensitive questions, we use prompt based temporal filtering(PBTF) module to better extract the features in temporal dimension. We first add a temporal position encoding to $\mathbf{V_{q}}$. Unlike the trainable position encoding used in the original Video-LLaMA, we design a position encoding based on the transformer’s[52] model. Experiments indicate that the revised position encoding is more effective. Furthermore, it facilitates the development of a multi-modal model with extrapolation capabilities in the future. The traditional transformer’s[52] position embedding is shown in Eq. 4, where pos represents the position, specifically the frame index here. i denotes the dimension, and d is the feature dimension. We improve upon it by substituting VPE for pos, where $VPE=pos\cdot(S/T)$, and $S$ is a constant. The ablation experiments demonstrate that our VPE enhances the effectiveness of STFM.
@@ 3.3.1 Zero-Shot Training
method PT Activitynet-QA MSVD-QA MSRTT-QA Next-QA
@@
Table 3: The few-shot evaluation results of various models on ActivityNet-QA, MSVD-QA, MSRTT-QA, and Next-QA. * indicates GPT-3.5-Turbo evaluation, while no * means non-GPT-3.5-Turbo evaluation. Bold represents the first place, and the underscore indicates the second place. PT refers to the number of pre-training datasets.
benchmark Activitynet-QA MSVD-QA MSRTT-QA Moviechat1k
@@
Video-LLaMA 12.4/1.1 51.6/2.5 29.6/1.8 51.7/2.57
FocusChat 33.2/3.1 52.4/3.4 46.7/3.2 60.0/3.54
difference value +20.8/+2.0 +0.8/+0.9 +17.1/+1.4 +8.3/+0.97
Table 4: The zero-shot evaluation results of FocusChat and Video-LLaMA on the ActivityNet-QA, MSVD-QA, MSRTT-QA, and MovieChat1K datasets, with the “difference value” representing the performance gap where FocusChat exceeds Video-LLaMA.
@@ 4.1 Implementation Details
We use the ViT-G/14 from EVA-CLIP[13] as the image encoder and CLIP[40] as the text encoder to ensure that the extracted text features and corresponding visual features are aligned in the embedding space. The image Q-Former and text Q-Former are initialized with the InstructBLIP’s[11] checkpoint, while the text Q-Former and video Q-Former are randomly initialized. We use the open-source LLaMA2 (7B) model as the LLM. In STFM, the parameters $\mathbf{\alpha}$ and $\mathbf{\beta}$ are set to 1 and 0, respectively, and the number of queries for all Q-Formers is 32. Hyper-parameters for all three zero-shot training stages are provided in Tab. 2. The few-shot training parameters are the same as those for zero-shot, except for the learning rate, which is set to 5e-5. Each few-shot training sample also includes the instruction: “Please answer as briefly as possible.”
@@ 4.2 Datasets
Our zero-shot model evaluation benchmarks include ActivityNet-QA [70], MSVD-QA[60], MSRTT-QA[60], and MovieChat1K[45], all of which are open-ended visual question-answering datasets. Except for MovieChat1K[45] with an average duration of 8 minutes, the video durations of the other datasets are around 1 to 2 minutes. For these evaluations, we use the widely adopted GPT-3.5-Turbo. For the few-shot experiments, we conduct training and testing on ActivityNet-QA[70], MSVD-QA[60], MSRTT-QA, and Next-QA[60]. Next-QA[58] has an average duration of 42 seconds and is a multiple-choice dataset. Since there is no standardized evaluation method for open-domain visual question-answering datasets in few-shot tasks, we used two evaluation approaches for all datasets except Next-QA[58]: one with GPT-3.5-Turbo and another with a strict evaluation method, where a prediction is considered correct only if it exactly matches the ground truth.
@@ 4.3 Main Results
Zero-Shot Result. For the quantitative experiments, we use a very small training dataset of approximately 1.5 million samples, which makes a direct comparison with models trained on tens of millions of unfair to some extent. Our model is an improvement on the Video-LLaMA structure. Thus, we mainly compared FocusChat with Video-LLaMA. Video-LLaMA uses a dataset of tens of millions of samples, whereas we use only about one-tenth of that amount. Despite this, FocusChat outperforms Video-LLaMA across all benchmarks.
@@
As shown in Tab. 4, on the ActivityNet-QA dataset, FocusChat’s accuracy exceeds Video-LLaMA by 20.8, with a score improvement of around 2. On the MSRTT-QA dataset, FocusChat’s accuracy surpasses Video-LLaMA by 17, with a score increase of 1.4. For MovieChat1K, we only test the global model, and FocusChat’s accuracy outperforms Video-LLaMA by 8.3, demonstrating FocusChat’s superior ability to understand long videos.On the MSVD-QA dataset, FocusChat matches Video-LLaMA in accuracy, with its score 0.94 higher.
FocusChat’s average score across all datasets is approximately 3.3, reflecting the high quality of its responses. FocusChat uses a smaller dataset compared with some other “SOTA” models, with model complexity nearly identical to that of Video-LLaMA. For the qualitative experiments, we compared the zero-shot performance of Video-LLaMA and FocusChat, as shown in Fig. 4, which demonstrates the accuracy of our model’s responses to different types of user prompts.
Few-Shot Result. To further validate the effectiveness of FocusChat, we conduct few-shot experiments based on a zero-shot model on open-domain visual question-answering datasets ActivityNet-QA, MSVD-QA, MSRTT-QA, and the multiple-choice dataset Next-QA. As shown in Tab. 3, we compare recent few-shot models on these datasets, many of which use quite large pre-training datasets. For instance, VideoCoCa achieves state-of-the-art performance on ActivityNet-QA with a pre-training dataset size as large as 4.8B, whereas our model only uses 0.72M pre-training data, still yields comparable results with SOTA performance. FocusChat gets 63.7 with GPT-3.5-Turbo evaluation on MSVD-QA, 54.6 without GPT-3.5-Turbo evaluation. The accuracy on Next-QA is 68.20, second only to the SeViLA model. This demonstrates the significant potential of FocusChat.

### v2

@@ 1 Introduction
Figure 3: When asking a question about a five-minute video, such as “How many people are playing basketball in the first two minutes?”, the semantic similarity matrix H and the attention-map SAM diagram in STFM are presented as indicative illustrations.
@@ 2 Related Work
Vision Large Language Models (vLLMs). The integration of the perceptual abilities of the vision models [21, 35, 37, 40] with the reasoning capabilities of LLMs has given rise to Vision Large Language Models (vLLMs) [66, 17, 27]. VLLMs encompass both Image-Language and Video-Language models. This approach transforms visual signals into tokens interpretable by LLMs. Image-Language models integrate powerful pre-trained language models with image encoders to enhance multi-modal reasoning capabilities[23, 30, 73]. For instance, Flamingo [1] connects state-of-the-art vision-only and language-only models to excel in few-shot learning tasks. BLIP-2[23] introduces a lightweight querying transformer to bridge the gap between frozen image encoders and language models, achieving strong performance with fewer trainable parameters. LLaVA [30] utilizes a simple linear layer to project image features into the text embedding space, effectively fine-tuning language models for improved outcomes. MiniGPT-4[73] builds on BLIP-2[23]by gathering a large dataset of image-text pairs, enhancing language generation. Video-language models have evolved from image-language models like Flamingo [1] and BLIP-2[23], which flatten spatio-temporal features into 1D sequences but struggle to capture temporal dynamics. Models such as Video-LLaMA[71] add video querying transformers to enhance temporal understanding, while Video-ChatGPT[33] averages spatial-level features for video representation. ChatVideo[53] uses tracklets annotated with textual descriptions and employs ChatGPT for various tasks, while VideoChat [24] generates action and object annotations for LLM reasoning.
@@
Long Video Understanding. Long video understanding poses significant challenges in computer vision, primarily because it requires capturing long-range patterns in videos that often exceed 30 seconds. A typical strategy[3, 55, 6] involves maintaining a memory bank to store historical information and utilizing parametric[55] or non-parametric[3] compression modules to manage this efficiently. Recent methodologies [19, 16, 41] have incorporated language as an intermediary for enhanced comprehension, segmenting extended videos into shorter clips, producing textual descriptions for each segment, and subsequently utilizing large language models (LLMs) to consolidate these captions for further analysis. However, these methods are cumbersome and lack conciseness. Some have adopted streaming methods[16, 39, 72, 42] to process long videos. Although these methods appear promising, indiscriminately handling all video frames inevitably leads to information redundancy. [47] is a key-frame-based model, distinct from previous methods. Although it appears to treat each frame differently, the key frame is fixed and unrelated to the prompt. Our proposed text-guided approach effectively addresses this issue. There are very few video understanding models based on this method, with [39] being one of them. However, it has several drawbacks, including a complex model structure, time-consuming computations, and an inability to filter information at the frame level or even at the region level. In contrast, our FocusChat effectively overcomes these limitations.
@@ 3.1.1 Visual Semantic Branch
The visual semantic branch extracts the semantic information from each image. It consists of a frozen pre-trained image encoder and a trainable image Q-Former. A video consists of $T$ frames is represented as $\mathbf{V}\in\mathbb{R}^{T\times H\times W\times 3}$. The input frames are passed into a pre-trained visual encoder, i.e., ViT, to obtain the visual frame features $\mathbf{V_{f}}\in\mathbb{R}^{T\times N_{v}\times C}$, where $N_{v}$ represents the number of patches, and $C$ denotes the number of feature channels. Subsequently, an image Q-Former further compresses the frame features. As Fig. 2 illustrates, the image Q-Former takes as input M learnable queries of length D. These queries interact with the frame features via cross-attention and update the initial queries to output the final M visual semantic vectors of length D, denoted as $\mathbf{V_{q}}\in\mathbb{R}^{T\times M\times D}$. Each visual vector contains semantic information at the region level or frame level. For example, the visual semantics extracted from the video in Fig. 1 may include elements such as men, women, vegetables, treadmills, etc.
@@ 3.1.2 Textual Semantic Branch
The text semantic branch is designed to semantically encode the user’s input prompt $P$. A pre-trained CLIP[40] model encodes the user prompt, resulting in prompt features $\mathbf{P_{f}}\in\mathbb{R}^{N_{p}\times D}$ , where $D$ denotes the dimension of text embedding, and $N_{p}$ is the text token number. We then employ a Q-Former, similar to the visual semantic branch, to encode $\mathbf{P_{f}}$. In this process, $\mathbf{P_{f}}$ is projected as keys and values. By utilizing M trainable query embeddings, we obtain M text vectors, denoted as $\mathbf{P_{q}}\in\mathbb{R}^{M\times D}$. Each textual vector contains semantic information at the word or subword level. For example, the prompt “How many people are cooking in the kitchen?” in Fig. 1 may be parsed into semantic vectors representing person, cooking, kitchen, etc.
@@ 3.2 Spatial-Temporal Filtering Module (STFM)
3.2.1 Prompt Based Temporal Filtering Module
As mentioned earlier, previous work lacks information filtering at the frame level. Therefore, we first propose a “Prompt-based Temporal Filtering Module” to assist our model in filtering information at the temporal level. In this module, we first add a temporal position encoding to $\mathbf{V_{q}}$. Unlike the trainable position encoding used in the original Video-LLaMA, we design a position encoding based on the transformer’s[52] model. Experiments indicate that the revised position encoding is more effective. The traditional transformer’s[52] position embedding is shown in Eq. 1, where pos represents the position, specifically the frame index here. i denotes the dimension, and d is the feature dimension. We improve upon it by substituting VPE for pos, where $VPE=pos\cdot(S/T)$, and $S$ is a constant, T is the number of input frames. The ablation experiments demonstrate that our VPE enhances the effectiveness of STFM.
The $\mathbf{V_{q}}$ added with temporal position encoding are then projected into keys and values using $\mathbf{W_{k}}$ and $\mathbf{W_{v}}$, respectively. Meanwhile, another set of N trainable queries is concatenated with the text semantic features $\mathbf{P_{q}}$ and used as input to a self-attention module to incorporate time-sensitive information into the queries. This helps our model extract visual features that aligned with the temporal information during hybrid spatial-temporal filtering cross-attention, enhancing the signal-to-noise ratio. For example, the visual features of the basketball scene in the last minute of Fig. 2 will also be filtered out as redundant information. Concretely , we only use the output of the self-attention corresponding to the trainable queries and linearly project it as the Query in Fig. 2, which is referred to as $\mathbf{V_{s}}\in\mathbb{R}^{N\times D}$ in the following text. Subsequently, $\mathbf{V_{s}}$ is multiplied by the keys, which contain temporal position encoding information, resulting in the prompt-based temporal feature. This prompt-based temporal feature, and the values, which also contain temporal position encoding information, will be fed into the next module to provide features that emphasize temporal filtering.
@@ 3.2.2 Hybrid Spatial-Temporal Filtering Module
In addition to the features that emphasize temporal filtering, we also need to obtain visual features that are semantically close to the user’s query, we first compute the semantic similarity matrix by calculating the similarity between text and visual semantic vectors. The similarity matrix is then used to guide the hybrid spatial-temporal filtering (HSTF) module for visual information filtering. The filtering process is explicitly applied to the video in both spatial and temporal dimensions. Spatial filtering is performed row by row, as shown in Fig. 3, while temporal filtering is applied column by column.
Semantic Similarity Matrix. We compute the cosine similarity between the text semantic feature $\mathbf{P_{q}}$ and the visual semantic feature $\mathbf{V_{q}}$ to get the similarity matrix $\mathbf{H}\in\mathbb{R}^{T\times M}$. As shown in Eq. 2, sim indicates cosine similarity and normalization, t is the frame index, and i is the semantic feature index. Since text semantics can achieve word-level granularity and visual semantics can achieve region-level granularity, therefore $\mathbf{H}$ could represents fine-grained similarity. For example, in Fig. 3, each element of the semantic similarity matrix H represents a specific region in the image, such as an arm, body, or basketball. It serves as a guide for filtering the visual spatial-temporal features.
@@
Hybrid Spatial-Temporal Filtering Cross-Attention. We aim to produce an attention map that is closely related to the semantic similarity matrix. Our hybrid spatial-temporal filtering module based attention-map $\mathbf{SAM}\in\mathbb{R}^{N\times TM}$ is given in Eq. 3, which is equivalent to $SAM=softmax(\alpha log(H)+QK^{T}/\sqrt{d})$, where $\alpha$ is a constant greater than or equal to 0, Q is the Query of PBTF, and K is the Key of PBTF. So $g(H)=\alpha log(H)$, as shown in Fig. 2. The output of the hybrid spatial-temporal filtering module is denoted as $\mathbf{Z}\in\mathbb{R}^{N\times d}$. Eq. 3 and Eq. 4 complete most of the spatiotemporal information filtering. Therefore, the redundant information filtered in $\mathbf{V_{q}}$ alleviates the pressure on the number of vision tokens sent to LLM.
@@ 3.3.1 Zero-Shot Training
method PT Activitynet-QA MSVD-QA MSRVTT-QA Next-QA
@@
Table 3: The few-shot evaluation results of various models on ActivityNet-QA, MSVD-QA, MSRVTT-QA, and Next-QA. * indicates GPT-3.5-turbo-0613 evaluation, while no * means non-GPT-3.5-turbo-0613 evaluation. Bold represents the first place, and the underscore indicates the second place. PT refers to the number of pre-training datasets.
benchmark Activitynet-QA MSVD-QA MSRVTT-QA MovieChat-1K
@@
Video-LLaMA 12.4/1.1 51.6/2.5 29.6/1.8 51.7/2.6
FocusChat 33.2/3.1 52.4/3.4 46.7/3.2 60.0/3.5
difference value +20.8/+2.0 +0.8/+0.9 +17.1/+1.4 +8.3/+0.9
Table 4: The zero-shot evaluation results of FocusChat and Video-LLaMA on the ActivityNet-QA, MSVD-QA, MSRVTT-QA, and MovieChat-1K datasets, with the “difference value” representing the performance gap where FocusChat exceeds Video-LLaMA.
@@ 4.1 Implementation Details
We use the ViT-G/14 from EVA-CLIP[13] as the image encoder and CLIP[40] as the text encoder to ensure that the extracted text features and corresponding visual features are aligned in the embedding space. The image Q-Former and text Q-Former are initialized with the InstructBLIP’s[11] checkpoint, while the STFM is randomly initialized. We use the open-source LLaMA2 (7B) model as the LLM. In STFM, the parameters $\mathbf{\alpha}$ and $\mathbf{\beta}$ are set to 1 and 0, respectively, and the number of queries for all Q-Formers is 32. Hyper-parameters for all two zero-shot training stages are provided in Tab. 2. The few-shot training parameters are the same as those for zero-shot, except for the learning rate, which is set to 5e-5. Each few-shot training sample also includes the instruction: “Please answer as briefly as possible.”
@@ 4.2 Datasets
Our zero-shot model evaluation benchmarks include ActivityNet-QA [70], MSVD-QA[60], MSRVTT-QA[60], and MovieChat-1K[45], all of which are open-ended visual question-answering datasets. Except for MovieChat-1K[45] with an average duration of 8 minutes, the video durations of the other datasets are around 1 to 2 minutes. For these evaluations, we use the widely adopted GPT-3.5-turbo-0613. For the few-shot experiments, we conduct training and testing on ActivityNet-QA[70], MSVD-QA[60], MSRVTT-QA, and Next-QA[60]. Next-QA[58] has an average duration of 42 seconds and is a multiple-choice dataset. Since there is no standardized evaluation method for open-domain visual question-answering datasets in few-shot tasks, we used two evaluation approaches for all datasets except Next-QA[58]: one with GPT-3.5-turbo-0613 and another with a strict evaluation method, where a prediction is considered correct only if it exactly matches the ground truth.
@@ 4.3 Main Results
Zero-Shot Result. For the quantitative experiments, we use a very small training dataset of approximately 1.5 million samples, which makes a direct comparison with models trained on tens of millions of samples somewhat unfair. Our model builds upon and enhances the Video-LLaMA structure to achieve better video understanding. Thus, we mainly compared FocusChat with Video-LLaMA. Video-LLaMA uses a dataset of tens of millions of samples, whereas we use only about one-tenth of that amount. Despite this, FocusChat outperforms Video-LLaMA across all benchmarks.
@@
As shown in Tab. 4, on the ActivityNet-QA dataset, FocusChat’s accuracy exceeds Video-LLaMA by 20.8, with a score improvement of around 2. On the MSRVTT-QA dataset, FocusChat’s accuracy surpasses Video-LLaMA by 17, with a score increase of 1.4. For MovieChat-1K, we only test the global mode, and FocusChat’s accuracy outperforms Video-LLaMA by 8.3, demonstrating FocusChat’s superior ability to understand long videos.On the MSVD-QA dataset, FocusChat matches Video-LLaMA in accuracy, with its score 0.94 higher.
FocusChat’s average score across all datasets is approximately 3.3, reflecting the high quality of its responses. For the qualitative experiments, we compared the zero-shot performance of Video-LLaMA and FocusChat, as shown in Fig. 4, which demonstrates the accuracy of our model’s responses to different types of user prompts.
Few-Shot Result. To further validate the effectiveness of FocusChat, we conduct few-shot experiments based on a zero-shot model on open-domain visual question-answering datasets ActivityNet-QA, MSVD-QA, MSRVTT-QA, and the multiple-choice dataset Next-QA. As shown in Tab. 3, we compare recent few-shot models on these datasets, many of which use quite large pre-training datasets. For instance, VideoCoCa achieves state-of-the-art performance on ActivityNet-QA with a pre-training dataset size as large as 4.8B, whereas our model only uses 0.72M pre-training data, still yields comparable results with SOTA performance. FocusChat gets 63.7 with GPT-3.5-turbo-0613 evaluation on MSVD-QA, 54.6 without GPT-3.5-turbo-0613 evaluation. The accuracy on Next-QA is 68.20, second only to the SeViLA model. This demonstrates the significant potential of FocusChat.